# Radio Map Prediction from Aerial Images and Application to Coverage Optimization

### v1

@@ VI-A Coverage Definition and Optimization Objective
As the first scenario, we consider a macro-diversity system [macro-diversity], in which signals from multiple BS are combined non-coherently and therefore the total received power in linear scale is defined as the sum
@@
$\begin{split}&\operatorname{SINR}(\vec{\varphi},\vec{\theta},x,y)\\ =&\max_{i=1,\ldots,M}\frac{\left(\operatorname{P}_{i}(\varphi_{i},\theta_{i},x,y)\right)_{W}}{\sum_{\begin{subarray}{c}j\neq i\end{subarray}}\left(\operatorname{P}_{j}(\varphi_{i},\theta_{i},x,y)\right)_{W}+(\operatorname{P}_{N})_{W}},\end{split}$
@@
VI-B Implementation and Results
@@
TABLE VI: Results of scenario 1 (4) in dBm
Minimizing (4) directly with gradient descent is problematic, since in each iteration only the value at one or a few specific locations with the current $10$-th percentile is taken into account during backpropagation. To obtain an approximate formulation, we replace the $10$-th percentile in (4) by the Boltzmann operator [boltzmann], defined as
@@
TABLE VII: Results of scenario 2 (5), number of pixels in the target area with SINR above the threshold
We consider $M=3$ BS, a $120^{\circ}$ sector antenna pattern, noise power of $-104$dBm (corresponding to a noise power spectral density of $-174$dBm/Hz and $10$MHz bandwidth), and we assume a Tx input power of $23$dBm as detailed in Section III-B. Both methods are run for $500$ iterations, which takes about $15$s with random search and $60$s with gradient descent, although we observed that gradient descent is already close to the optimal value after about $50$ iterations. The SINR threshold in (5) is set to $t=20\operatorname{dB}$.
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 7: Coverage Optimization
The results are listed in Tables VI and VII and visualized in Figures 7(a) and 7(b). Both methods provide a solid improvement over the baseline in scenario 1 and a large improvement in scenario 2, with gradient descent performing slightly better. This also shows that our approach of optimizing a differentiable approximation of the coverage score with gradient descent is indeed valid in order to maximize the true score.
VII CONCLUSION
@@
As a further potential extension of this, predicting the radio map from satellite data available for the whole planet (similarly to [plgan], but in cellular networks) would allow large-scale applications to network planning and related tasks. The often lower spatial resolution may make the recognition of objects and shapes more difficult. On the other hand, incorporation of other spectral bands or radar data could be beneficial for the implicit classification of objects and height estimation [sentinel_height].
Although this study does not cover it, the dataset could be used to explore joint semantic segmentation and height estimation from aerial images [im2height], [joint_seg_height]. This could serve as an intermediate step toward improving radio map predictions.
By making our dataset and code available for public use, we aim to facilitate the work of other researchers and promote reproducible and comparable investigations. Featuring directive Tx antennas, the dataset opens potential for the investigation of more downstream tasks in 5G/6G networks, such as beam codebook design and beam management [beams].

### v2

@@ IV-A Coverage Definition and Optimization Objective
As the first scenario, we consider a macro-diversity system [39], in which signals from multiple BS are combined non-coherently and therefore the total received power in linear scale is defined as the sum
@@
$\begin{split}&\operatorname{SINR}(\vec{\varphi},\vec{\theta},x,y)\\ =&\max_{i=1,\ldots,M}\frac{\left(\operatorname{P}_{i}(\varphi_{i},\theta_{i},x,y)\right)_{W}}{\sum_{\begin{subarray}{c}j\neq i\end{subarray}}\left(\operatorname{P}_{j}(\varphi_{i},\theta_{i},x,y)\right)_{W}+(\operatorname{P}_{N})_{W}},\end{split}$ (5)
@@
IV-B Implementation and Results
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
(${\mapsto}$ Refer to caption Refer to caption
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
(${\mapsto}$ Refer to caption Refer to caption
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
(${\mapsto}$ Refer to caption Refer to caption
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
(${\mapsto}$ Refer to caption Refer to caption
Figure 9: Coverage Optimization
@@ IV-B Implementation and Results
TABLE VII: Results of scenario 1 (4) in dBm
Minimizing (4) directly with gradient descent is problematic, since in each iteration only the value at one or a few specific locations with the current $10$-th percentile is taken into account during backpropagation. To obtain an approximate formulation, we replace the $10$-th percentile in (4) by the Boltzmann operator [40], defined as
@@
TABLE VIII: Results of scenario 2 (6), number of pixels in the target area with SINR above the threshold
We consider $M=3$ BS, a $120^{\circ}$ sector antenna pattern, noise power of $-104$dBm (corresponding to a noise power spectral density of $-174$dBm/Hz and $10$MHz bandwidth), and we assume a Tx input power of $23$dBm as detailed in Section II-B2. Both methods are run for $500$ iterations, which takes about $15$s with random search and $60$s with gradient descent, although we observed that gradient descent is already close to the optimal value after about $50$ iterations. The SINR threshold in (6) is set to $t=20\operatorname{dB}$.
The results are listed in Tables VII and VIII and visualized in Figures 9(a) and 9(b). Both methods provide a solid improvement over the baseline in scenario 1 and a large improvement in scenario 2, with gradient descent performing slightly better. This also shows that our approach of optimizing a differentiable approximation of the coverage score with gradient descent is indeed valid in order to maximize the true score.
V CONCLUSION
@@
As a further potential extension of this, predicting the radio map from satellite data available for the whole planet (similarly to [22], but in cellular networks) would allow large-scale applications to network planning and related tasks. The often lower spatial resolution may make the recognition of objects and shapes more difficult. On the other hand, incorporation of other spectral bands or radar data could be beneficial for the implicit classification of objects and height estimation [41].
In this work, we made certain assumptions, such as approximating the ground as flat, to simplify the modeling process. We acknowledge that these simplifications may limit the applicability of our models to regions with significant elevation variations. Future work could involve retraining models with datasets that include ground elevation data to better accommodate diverse geographic scenarios, as in [15]. While we have presented the radiomap prediction results in terms of the RMSE averaged over the whole map, in line with the existing literature, a more refined study of the “conditional” RMSE (e.g., RMSE versus distance from Tx, or RMSE conditioned on (non-)line-of-sight propagation) is certainly an interesting topic for future work.
Although this study does not cover it, the dataset could be used to explore joint semantic segmentation and height estimation from aerial images [38], [42]. This could serve as an intermediate step toward improving radio map predictions.
By making our dataset and code available for public use, we aim to facilitate the work of other researchers and promote reproducible and comparable investigations. Featuring directive Tx antennas, the dataset opens potential for the investigation of more downstream tasks in 5G/6G networks, such as beam codebook design and beam management [3].